# LEASE: Offline Preference-based Reinforcement Learning with High Sample Efficiency

### v1

@@ VII-D Results for Policy Performance
The effect of selecting mechanism for agent performance. Table VI compares the agent performance between the proposed method LEASE and the method not using selecting mechanism (We denote this method as FRESH). This table indicates that the application of selecting mechanism $f(\sigma_{0},\sigma_{1})$ can efficiently improve agent performance. This results also further validates the effectiveness and advantages of the designed selecting mechanism.
@@
The effect of the number of preference data $N_{l}$ for agent performance. Fig. 4 indicates that the number of preference data $N_{l}$ influences the accuracy of reward model directly. URLHF [21] trains reward model with $2000$ preference dataset. Our method LEASE aims to achieve comparable performance with URLHF under fewer preference dataset. Table IX shows the agent performance when the number of preference dataset is $20$ and $100$, where we use CQL algorithm for mujoco tasks and IQL algorithm for adroit tasks. This table indicates LEASE still can improve agent performance under very fewer preference dataset. Notably, the performance gap between $100$ preference data and $20$ preference data is not much. This indicates that LEASE have potential to perform well under more fewer preference dataset.
VIII Discussion
VIII-A Discussion for Offline PbRL Theory
@@
In offline RL field, high sample efficiency refers that the agent can achieve comparable performance under fewer data compared with the performance under large data. In this paper, the data refers to the preference dataset. The labeled preference dataset, each trajectory of length $L$, is collected through real-time human feedback under policy $\mu$, which demands tremendous human effort, thus the collected cost of preference data is higher than fixed offline data. The unlabeled dateset is generated through trained transition without real-time interaction under learned policy $\pi^{t}$ at time $t$.
@@
Through data augmentation, the sample efficiency can be significantly reduced. The performance gap caused by reward can be reduced to $\varpi$ under fewer labeled dataset. Zhan et al.[18] also developed systematical theory for offline PbRL and introduced the concentrability coefficient $\mathscr{C}_{r}(\pi)$ for PbRL, defined as
@@
where $r(\sigma^{0},\sigma^{1})=\sum_{i}[R(s_{i}^{0},a_{i}^{0})-R(s_{i}^{1},a_{i}^{1})]$, $\pi_{\text{ref}}$ is an arbitrary trajectory distribution (usually set as $\mu_{1}$), and $\mu_{0},\mu_{1}$ are behavior trajectory distribution. However, it is based on trajectory and is difficult to combine with other offline RL theories. The proposed concentrability coefficient in our paper defined in Eq. (13) is based on state-action pairs.
Moreover, the performance gap of offline PbRL between behavior policy and learned policy is influenced by offline algorithm itself and the performance of the learned reward (preference) model. However, [18] fails to consider the gap caused by offline algorithm itself and the sampling error. The theory developed in our paper can be easily combined with other offline algorithm. The method in [18] can learn $\varpi$-optimal policy with a sample complexity of
@@
where $c>0$ is a universal constant, $k=(\inf_{x\in[-r_{max},r_{max}]}\\ \Phi^{\prime}(x))^{-1}$, and $\Phi(x)$ is a monotonically increasing link function. Compared with Eq. (17), the sample complexity of LEASE defined in Eq. (15) contains more useful information. For example, the sample complexity is related to the length of preference data and the range of reward function, and can be reduced when the learned policy is closed to behavior policy.
VIII-B Discussion for Broader Impacts
@@
Actual application. In actual scenarios, especially for human-in-loop-control, such as exoskeleton robot assistance or rehabilitation, designing a high sample efficient RL algorithm is greatly significant. Firstly, many rewards are difficult to describe in mathematical terms in some situation, such as the comfort of interaction. Human feedback or preference is the better way to reflect the above indexes. In addition, in human in-the-loop control, obtaining preference data through interaction can easily cause fatigue, and inappropriate interaction may cause damage to human. It is a necessity to learn reward model from limited preference dataset. Therefore, PbRL has a great potential to improve control performance in some scenarios where the reward function is difficult to describe and human is in the control loop.
@@
Theoretical study. The theory of LEASE shows that the performance gap between the behavior policy $\mu$ and $\widehat{\pi}$ learned by LEASE includes two part: the gap $\xi$ caused by offline algorithm itself and the gap $\xi_{1}$ caused by reward model gap (the details see Theorem 2), where the term $\xi_{1}$ only depends on preference dataset. Moreover, the theory of LEASE is based on state-action pairs, which is consistent with most offline algorithms theory. Therefore, the theory developed in this paper can be easily combined with other offline algorithms theory and easily used to build theory of policy improvement guarantee, which can provide the theoretical basis for offline PbRL and facilitate the development of offline PbRL theory.
IX Conclusion
@@
This paper proposes a novel offline PbRL algorithm (LEASE) with high sample efficiency. LEASE can achieve comparable performance under fewer preference dataset. By selecting high confidence and low variance data, the stability and accuracy of reward model are guaranteed. Moreover, this paper provides the theoretical analysis for LEASE, including generalization of reward model and policy improvement guarantee. This theory can be easily connected with other offline algorithms. The theoretical and experimental results demonstrate that the data selecting mechanism $f(\sigma_{0},\sigma_{1})$ can effectively improve performance of reward model and the performance learned by LEASE can be guaranteed under fewer preference dataset. However, there is still the gap between the true and the learned reward model. Future works can focus on how to further reduce this gap under the limited preference data or how to achieve conservative estimation for state-action pairs where the learned reward model predicts inaccurately.
@@ A-A Proof of Proposition 1
$\eta_{1}=\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5,y^{j}=0\Big{]},$ (A.1)
$\eta_{2}=\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5,y^{j}=1\Big{]},$
where $\eta_{1}$ represents the error ratio of classifying the first category as the second category, and $\eta_{2}$ is the error ratio of classifying the second category as the first category. Then, we prove the gap between the unlabeled loss with pseudo-label ^$\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ from two sides: ^$\widehat{\mathcal{L}}_{u}(\psi)\leq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega$ and ^$\widehat{\mathcal{L}}_{u}(\psi)\geq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega$. Notably, we ignore the selecting mechanism in the below proof since it doesn’t influence the proof result.
Step 1: Proving upper bound : ^$\widehat{\mathcal{L}}_{u}(\psi)\leq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega$.
^$\widehat{\mathcal{L}}_{u}(\psi)$
@@
$=$ $-\frac{1}{{N}_{u}}\sum_{j=1}^{{N}_{u}}\mathbbm{1}\big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5\big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
$+\mathbbm{1}\big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5\big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
@@
$+\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5,y^{j}=0\Big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
$+\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5,y^{j}=1\Big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
$\leq$ ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta_{1}\max_{j}\Big{\{}-\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}\Big{\}}$
@@
$\leq$ ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega.$
Step 2: Proving low bound : ^$\widehat{\mathcal{L}}_{u}(\psi)\geq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega$.
^$\widehat{\mathcal{L}}_{u}(\psi)$
@@
$-\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{1}^{j}\succ\sigma_{0}^{j}\right)>0.5,y^{j}=0\Big{]}\log P_{\psi}\big{(}\sigma_{0}^{j}\succ\sigma_{1}^{j}\big{)}$
$-\mathbbm{1}\Big{[}P_{\psi}\left(\sigma_{0}^{j}\succ\sigma_{1}^{j}\right)>0.5,y^{j}=1\Big{]}\log P_{\psi}\big{(}\sigma_{1}^{j}\succ\sigma_{0}^{j}\big{)}$
$\geq$ ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega.$
Combing Step 1 and Step 2, we can obtain the following result:
|^$\Big{|}\widehat{\mathcal{L}}_{u}(\psi)-\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\Big{|}\leq\eta\Omega.$ (A.2)
@@ A-B Proof of Theorem 1
where $\mathbb{E}_{\widehat{\mathcal{D}}}\left[f\right]=\sum_{i=1}^{m}\left[g(x_{i})\right]/m$ is the empirical form of $\mathbb{E}_{\mathcal{D}}\left[f\right]$. This proof can be found in Theorem 3.3 of work [24].
@@
^$\widehat{\Re}_{\widehat{\mathcal{D}}}(\mathcal{F})\leq 2\widehat{\Re}_{\widehat{\mathcal{D}}}\big{(}\Pi(\mathcal{R})\big{)}.$ (A.4)
@@
The Proof for Theorem 1: This proof is based on Proposition 1 and 2. Combing Eqs. (A.3) and (A.4), we can derive that for labeled dataset $\widehat{\mathcal{D}}=\{(\sigma_{0},\sigma_{1},y)^{(i)}\}_{i=1}^{N}$, the generalization error bound between expected error $\mathcal{L}(\psi)$ and empirical error ^$\widehat{\mathcal{L}}(\psi)$ holds:
@@
However, the dataset for training reward model includes two parts: the labeled dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$ and unlabeled dataset $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$, and empirical error ^$\widehat{\mathcal{L}}_{R}(\psi)=\widehat{\mathcal{L}}_{l}(\psi)+\widehat{\mathcal{L}}_{u}(\psi)$. Let ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ be the empirical error under unlabeled dataset with true label, then
@@
where $\widehat{\mathcal{D}}$ is the input combination of labeled and unlabeled dataset, denoted as $\{(\sigma_{0},\sigma_{1})^{(i)}\}_{i=1}^{N_{l}+N_{u}}$. According to Proposition 1, that is ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\leq\widehat{\mathcal{L}}_{u}(\psi)+\eta\Omega.$, we can derive

### v2

@@ VI-C2 The effect of selection mechanism on agent performance
To answer question $\mathbf{Q_{3}}$, we compare the agent performance between the proposed method LEASE and the method not using the selection mechanism (We denote this method as FRESH). Fig. 3 shows that the performance of the reward model can be significantly improved by using the selection mechanism. Table IV compares the results between LEASE and FRESH under three locomotion tasks based on CQL and IQL offline RL algorithms. This table indicates that the application of the selection mechanism $f(\sigma_{0},\sigma_{1})$ can efficiently improve agent performance, and the average performance of the policy learned by LEASE can be improved by $5.6\%$ compared with not using the selection mechanism. These results also further validate the effectiveness and advantages of the designed selection mechanism.
VI-C3 The effect of the number of preference data $N_{l}$ for agent performance
@@
Fig. 4 indicates that the number of preference data $N_{l}$ influences the accuracy of the reward model directly. URLHF [21] trains reward model with $2000$ preference dataset. This paper aims to achieve competitive performance with limited preference data, that is, achieve comparable performance with URLHF under a smaller preference dataset. We further investigate the impact of the number of preference data on the agent’s performance. Table V compares the agent’s performance when the number of preference datasets is $20$ and $100$, using the CQL algorithm for MuJoCo tasks and the IQL algorithm for Adroit tasks. This table indicates LEASE can still improve agent performance under a very small preference dataset. Notably, the performance gap between $100$ preference data and $20$ preference data is not much. This indicates that LEASE has the potential to perform well under a smaller preference dataset.
VI-D Parameter Analysis
TABLE VI: The results of parameter analysis for the confidence threshold $\kappa_{p}$ and the uncertainty threshold $\kappa_{\tau}$.
CQL $\kappa_{\tau}=0.05$ $\kappa_{\tau}=0.08$
$\kappa_{p}=0.85$ $\kappa_{p}=0.95$ $\kappa_{p}=0.85$ $\kappa_{p}=0.95$
walker2d-m $\mathbf{78.4\pm 0.9}$ $70.6\pm 1.8$ $72.4\pm 1.1$ $70.3\pm 2.1$
pen-expert $127.0\pm 3.4$ $\mathbf{130.6\pm 1.8}$ $122.9\pm 1.8$ $125.8\pm 2.4$
IQL $\kappa_{\tau}=0.05$ $\kappa_{\tau}=0.08$
$\kappa_{p}=0.85$ $\kappa_{p}=0.95$ $\kappa_{p}=0.85$ $\kappa_{p}=0.95$
walker2d-m $72.6\pm 0.1$ $69.1\pm 1.5$ $\mathbf{73.0\pm 2.1}$ $68.9\pm 3.3$
[…]
The selection mechanism $f(\sigma_{0},\sigma_{1})$ for the reward model has two critical parameters: the confidence threshold $\kappa_{p}$ and the uncertainty threshold $\kappa_{\tau}$. The selection of two parameters depends on multiple factors, including the complexity of the task’s reward function, the performance of the pre-trained reward model, and the specific offline RL algorithm employed (since the collected preference data is inherently linked to the policy learned by the algorithm). Given the substantial computational cost inherent in RL training, exhaustive parameter tuning is often impractical. To mitigate this, our work systematically analyzes the impact of key parameters on task performance across multiple algorithms, offering practical guidance for their configuration.
Table VI presents a comparative analysis of the performance of different offline RL algorithms (CQL and IQL) under varying parameter configurations for both locomotion and manipulation tasks. For the uncertainty threshold $\kappa_{\tau}$, the results demonstrate that CQL exhibits a smaller uncertainty threshold compared to IQL. This phenomenon can be attributed to CQL’s inherent conservatism, which leads to a narrower data distribution when the learned policy interacts with the environmental model (i.e., the generated data is more concentrated within a specific region). Consequently, the reward model exhibits lower prediction uncertainty, allowing for the selection of a reduced uncertainty threshold. For the confidence threshold $\kappa_{p}$, manipulation tasks require a higher confidence threshold than locomotion tasks. This discrepancy arises from the simpler reward structure of manipulation tasks, which typically rely on discrete variables (See Fig. A.1 in Appendix D). As a result, the pre-trained reward model achieves higher accuracy in manipulation tasks, allowing for the use of a more stringent confidence threshold to enhance the quality of generated data. For additional experiments of the LEASE—such as evaluations of the reward model on alternative tasks, its application within model-based RL algorithms, and an extended comparative analysis—please refer to Appendix D.
VII Discussion
VII-A Discussion for Offline PbRL Theory
@@
In the offline RL field, high sample efficiency refers to the agent’s ability to achieve comparable performance under fewer data compared with the performance under large data. In this paper, the data refers to the preference dataset. The labeled preference dataset, each trajectory of length $L$, is collected through real-time human feedback under policy $\mu$, which demands tremendous human effort; thus, the collected cost of preference data is higher than fixed offline data. The unlabeled dataset is generated through a trained transition without real-time interaction under the learned policy $\pi^{t}$ at time $t$.
@@
Through data augmentation, the sample efficiency can be significantly reduced. The performance gap caused by reward can be reduced to $\varpi$ under a smaller labeled dataset. Zhan et al.[18] also developed a systematic theory for offline PbRL, but did not focus on the generalization performance of the reward model. Moreover, they also introduced the concentrability coefficient $\mathscr{C}_{r}(\pi)$ for PbRL, defined as
@@
where $r(\sigma^{0},\sigma^{1})=\sum_{i}[R(s_{i}^{0},a_{i}^{0})-R(s_{i}^{1},a_{i}^{1})]$, $\pi_{\text{ref}}$ is an arbitrary trajectory distribution (usually set as $\mu_{1}$), and $\mu_{0},\mu_{1}$ are behavior trajectory distribution. However, it is based on trajectory and is difficult to combine with other offline RL theories. The proposed concentrability coefficient in our paper, defined in Eq. (13), is based on state-action pairs.
Moreover, the performance gap of offline PbRL between the behavior policy and learned policy is influenced by the offline algorithm itself and the performance of the learned reward (preference) model. However, [18] fails to consider the gap caused by the offline algorithm itself and the sampling error. The theory developed in our paper can be easily combined with another offline algorithm. Then, the method in [18] can learn $\varpi$-optimal policy with a sample complexity of
@@
where $c>0$ is a universal constant, $k=(\inf_{x\in[-r_{max},r_{max}]}\\ \Phi^{\prime}(x))^{-1}$, and $\Phi(x)$ is a monotonically increasing link function. Compared with Eq. (17), the sample complexity of LEASE defined in Eq. (15) contains more useful information. For example, the sample complexity is related to the length $L$ of preference data and the range $R_{max}$ of the reward function, and can be reduced when the learned policy is close to the behavior policy $\mu$.
VII-B Discussion for Broader Impacts
@@
Actual application. In actual scenarios, especially for human-in-loop control, such as exoskeleton robot assistance or rehabilitation, designing a highly sample-efficient RL algorithm is greatly significant. Firstly, many rewards are difficult to describe in mathematical terms in some situations, such as the comfort of interaction. Human feedback or preference is a better way to reflect the above indices. In addition, in human-in-the-loop control, obtaining preference data through interaction can easily cause fatigue, and inappropriate interaction may cause damage to humans. It is a necessity to learn a reward model from a limited preference dataset. Then, for robot assistance or rehabilitation, the goal of the robot is to reduce human metabolic cost and guarantee the comfort of interaction. However, the reward model for this goal is difficult to design. Therefore, PbRL has a great potential to improve control performance in some scenarios where the reward function is difficult to describe and humans are in the control loop.
@@
Theoretical study. The theory of LEASE shows that the performance gap between the behavior policy $\mu$ and $\widehat{\pi}$ learned by LEASE includes two part: the gap $\xi$ caused by offline algorithm itself and the gap $\xi_{1}$ caused by reward model gap (the details see Theorem 2), where the term $\xi_{1}$ only depends on preference dataset. Moreover, the theory of LEASE is based on state-action pairs, which is consistent with most offline algorithms theory. Therefore, the theory developed in this paper can be easily combined with other offline algorithm theories and easily used to build the theory of policy improvement guarantee, which can provide the theoretical basis for offline PbRL and facilitate the development of offline PbRL theory.
VIII Conclusion
@@
This paper proposes a novel offline PbRL algorithm (LEASE) with high sample efficiency. LEASE can achieve comparable performance under fewer preference datasets. By selecting high-confidence and low-variance data, the stability and accuracy of the reward model are guaranteed. Moreover, this paper provides the theoretical analysis for LEASE, including the generalization of the reward model and policy improvement guarantee. The theoretical and experimental results demonstrate that the data selection mechanism $f(\sigma_{0},\sigma_{1})$ can effectively improve the performance of the reward model, and the performance learned by LEASE can be guaranteed under fewer preference datasets. Future work can focus on reducing this gap further under limited preference data or on achieving conservative estimation for state-action pairs where the learned reward model predicts inaccurately.
@@ A-A Proof of Proposition 1
[…]
where $\eta_{1}$ represents the error ratio of classifying the first category as the second category, and $\eta_{2}$ is the error ratio of classifying the second category as the first category. Then, we prove the gap between the unlabeled loss with pseudo-label $\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ from two sides: $\widehat{\mathcal{L}}_{u}(\psi)\leq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega$ and $\widehat{\mathcal{L}}_{u}(\psi)\geq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega$. Notably, we ignore the selection mechanism in the proof below since it doesn’t influence the proof result.
Step 1: Proving upper bound : $\widehat{\mathcal{L}}_{u}(\psi)\leq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega$.
$\widehat{\mathcal{L}}_{u}(\psi)$
@@
[…]
@@
[…]
$\leq$ $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta_{1}\max_{j}\Big\{-\log P_{\psi}\big(\sigma_{0}^{j}\succ\sigma_{1}^{j}\big)\Big\}$
@@
$\leq$ $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)+\eta\Omega.$
Step 2: Proving low bound : $\widehat{\mathcal{L}}_{u}(\psi)\geq\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega$.
$\widehat{\mathcal{L}}_{u}(\psi)$
@@
[…]
$\geq$ $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)-\eta\Omega.$
Combining Step 1 and Step 2, we can obtain the following result:
$\Big|\widehat{\mathcal{L}}_{u}(\psi)-\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\Big|\leq\eta\Omega.$ (A.2)
@@ A-B Proof of Theorem 1
where $\mathbb{E}_{\widehat{\mathcal{D}}}\left[f\right]=\sum_{i=1}^{m}\left[g(x_{i})\right]/m$ is the empirical form of $\mathbb{E}_{\mathcal{D}}\left[f\right]$. This proof can be found in Theorem 3.3 of the work [24].
@@
$\widehat{\Re}_{\widehat{\mathcal{D}}}(\mathcal{F})\leq 2\widehat{\Re}_{\widehat{\mathcal{D}}}\big(\Pi(\mathcal{R})\big).$ (A.4)
@@
The Proof for Theorem 1: This proof is based on Proposition 1 and 2. Combing Eqs. (A.3) and (A.4), we can derive that for labeled dataset $\widehat{\mathcal{D}}=\{(\sigma_{0},\sigma_{1},y)^{(i)}\}_{i=1}^{N}$, the generalization error bound between expected error $\mathcal{L}(\psi)$ and empirical error $\widehat{\mathcal{L}}(\psi)$ holds:
@@
However, the dataset for training reward model includes two parts: the labeled dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$ and unlabeled dataset $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$, and empirical error $\widehat{\mathcal{L}}_{R}(\psi)=\widehat{\mathcal{L}}_{l}(\psi)+\widehat{\mathcal{L}}_{u}(\psi)$. Let $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ be the empirical error under an unlabeled dataset with true label, then
@@
where $\widehat{\mathcal{D}}$ is the input combination of labeled and unlabeled dataset, denoted as $\{(\sigma_{0},\sigma_{1})^{(i)}\}_{i=1}^{N_{l}+N_{u}}$. According to Proposition 1, that is $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\leq\widehat{\mathcal{L}}_{u}(\psi)+\eta\Omega.$, we can derive